# Tropicalization through the lens of Łukasiewicz logic, with a topos theoretic perspective

### v1

@@ 6 MV-algebras and the topos $\widehat{\mathbb{N}^{\times}}$
$G\subseteq\mathbb{Q}$ if and only if $G$ is a group, $G$ is torsion free, and for every $x,y\not=0\in G$ there are $m,n\in\mathbb{N}$ such that $mx=ny$. Furthermore, denoting the order of $G$ by $<_{G}$, then the following conditions hold:
$\forall x$ $\forall y$ $\exists n$ $\exists m$ such that: $m,n\in Z$
@@
The class $Perf^{\mathbb{Q}}$ is geometric in the language of $MV$-algebras..
@@
$\forall x$ $\forall y$ $\exists n$ $\exists m$ such that: $m,n\in Z^{+}$
@@
The previous results give information on the theory of certain subgroups of $\mathbb{Q}$. By using the functor $\Delta$ one can transfer this information on perfect MV-algebras.
Theorem 6.6.
The first order theory $T_{\mathbb{Q}}^{\Delta}$ of $Perf^{\mathbb{Q}}$ is mutually many one reducible with $T_{\mathbb{Q}}$. In particular it is recursively axiomatizable.
Proof.
The theorem follows from Theorem 6.2 and Theorem 5.2. ∎
The Theorem 6.2 has some corollaries.
Corollary 6.2.
Every two perfect MV-chains with divisible radical are elementarily equivalent.
Proof.
It is well known that every two divisible totally ordered abelian groups are elementarily equivalent. Moreover $G$ is divisible if and only if $\Delta(G)$ has divisible radical. Now the result follows from Theorem 6.2. ∎
Corollary 6.3.
The first order theory of perfect MV-chains is decidable.
Proof.
The theory of perfect MV-chains reduces to the theory of totally ordered abelian groups, which is decidable by [23]. ∎
The following result is a kind of converse of theorem 6.2, based on [10] prop 2.4.2.
We go from an abelian $\ell$-group $G$ to $G^{+}$.
Lemma 6.3.
Let $\phi$ be a formula in the language of lattice ordered groups. There is a formula $\Delta^{-1}+(\phi)$ in the language of lattice ordered monoids such that $G\models\phi(x_{1},\ldots,x_{n})$ if and only if $G^{+}\models\Delta^{-1}+(\phi)(x_{1}^{-},x_{1}^{+},\ldots,x_{n}^{-},x_{n}^{+}).$ Recall $x^{+}=x\vee 0$ and $x^{-}=x\wedge 0$.
Proof.
$x_{i}$ is replaced by $x_{i}^{-}-x_{i}^{+}$, and $\exists y.\psi(y)$ is replaced with $\exists z,w\in G^{+}.\psi(z-w)$. $\ell$-group Terms are inductively replaced by differences of lattice monoid terms as follows (see [10] prop 2.4.2 on how to construct the Grothendieck group of a lattice ordered monoid).
  • $0=0-$ 0 - 0;
  • $-(a-b)=b-a$;
  • $(a-b)+(c-d)=(a+c)-(b+d)$;
  • $(a-b)\wedge(c-d)=((a+d)\wedge(b+c))-(b+d)$;
  • $(a-b)\vee(c-d)=((a+d)\vee(b+c))-(b+d)$.
At the end every equality of $\ell$ group terms $t-u=v-w$ is replaced by $t+w=u+v$ which is a formula of lattice ordered monoids.
Now the lattice ordered monoids $(G^{+},+,\wedge,\vee)$ and $(Rad(\Delta(G)),\oplus,\wedge,\vee)$ are isomorphic. So we have:
Corollary 6.4.
Let $\phi$ be a formula in the language of lattice ordered groups. There is a formula $\Delta^{-1}(\phi)$ in the language of MV-algebras such that $G\models\phi(x_{1},\ldots,x_{n})$ if and only if $\Delta^{-1}(G)\models\Delta^{-1}(\phi)(x_{1}^{-},x_{1}^{+},\ldots,x_{n}^{-},x_{n}^{+}).$ Recall $x^{+}=x\vee 0$ and $x^{-}=x\wedge 0$.
We obtain the following model theoretic information:
Corollary 6.5.
The theory of perfect MV-algebras is undecidable.
Proof.
This follows by the previous lemma since the theory of all abelian $\ell$-groups is undecidable, see [5]. ∎
Finally we perform translations between first order formulas but we do not introduce any quantifier. So we can use these translations to prove model theoretic results like elimination of quantifiers or model completeness for theories of perfect MV-algebras.
We begin by observing:
Lemma 6.4.
The formula $\Delta(\phi)$ has the same quantifier depth as $\phi$ and , if $\phi$ is a positive boolean combination of atomic formulas, then so is $\Delta(\phi)$.
The same holds for the formula $\Delta^{-1}(\phi)$.
This gives, for instance, the following results: In terms of classical model theory we obtain:
Theorem 6.7.
The theory of perfect MV-chains with divisible radical is the model completion of the theory of perfect mv chains.
Proof.
This follows from the theorem of [28] that the theory of divisible MV chains is the model completion of the theory of mv chains. ∎
Theorem 6.8.
The theory of perfect MV-chains with divisible radical verifies the elimination of quantifiers. In particular, this theory is model complete, and every element of $Perf^{\mathbb{Q}}$ is an elementary substructure of $\Delta(\mathbb{Q})$.
Proof.
The theory T of divisible totally ordered abelian groups verifies the elimination of quantifiers. The perfect MV-algebra $\Delta(G)$ has a divisible radical if and only if $G$ is a divisible totally ordered abelian group. In $T$, every formula is equivalent to a quantifier free formula. By the previous subsection, the same holds for the theory of perfect MV-chains with divisible radical. ∎
@@
and $\Theta_{pt}$ is a functor from $Pt(\widehat{\mathbb{N}^{\times}})$ to $C\ell BS$.
@@
There is a natural functor from the category $Pt(\widehat{\mathbb{N}^{\times}})$ to the category Coupled $\ell$-Bisemirings.
@@
Indeed, we have a threefold interaction among the geometric side, points of the topos $\widehat{\mathbb{N}^{\times}}$, the algebraic side of ordered subgroups of $\mathbb{Q}$ and the logical side of $Perf^{\mathbb{Q}}$, models of an extension of $\L$ukasiewicz logic.
Given the categorial equivalence between $Perf^{\mathbb{Q}}$ and $Pt(\widehat{\mathbb{N}^{\times}})$, we derive:
Corollary 6.8.
The first order theory of the points of the topos $\widehat{\mathbb{N}^{\times}}$ is recursively axiomatizable.
Proof.
A point $\mathbf{p}\in\widehat{\mathbb{N}^{\times}}$ is a set $X$ together with a $\mathbb{N}$-indexed family of functions $f_{n}$ from $X$ to $X$. Suppose $\sigma$ is a sentence of the language of $Pt(\widehat{\mathbb{N}^{\times}})$, and $\sigma$ is true in a point $\mathbf{p}\in Pt(\widehat{\mathbb{N}^{\times}})$ where $X$ is a subgroup $H$ of $\mathbb{Q}$ and $f_{n}(x)=nx$. Then $\sigma$ is true in some subgroup of $\mathbb{Q}$. Likewise, if $\sigma$ is true in a subgroup of $\mathbb{Q}$, then $\sigma$ is true in point of $\widehat{\mathbb{N}^{\times}}$.
So, the first order theory of $Pt(\widehat{\mathbb{N}^{\times}})$ is the same of that of the subgroups of $\mathbb{Q}$ and is recursively axiomatizable.
@@ 6.1 Flatness and $Pt(\widehat{\mathbb{N}^{\times}})$
    $F(*)=X\not=\emptyset$, where * is the unique object of $\mathbb{N}^{\times}$
@@ 7 Conclusions
As stated in the Introduction, the purpose of this article is to illustrate the interactions between algebraic structures such as Semirings, Tropical structures and $\L$ukasiewicz logic.
This perspective has led to a generalisation of the concept of Tropicalisation of algebraic structures in a form that opened up an unexpected interpretation of an extension of $\L$ukasiewicz logic. This logic exhibits models that are functorially connected to a category of points of a non-commutative geometry.
Thus, assuming the functorial version of the tropicalization of an algebraic structure, the above interation, among others, takes form as follows:
  • from the algebraic side arise the new algebraic categories of Lattice Bisemirings and of Coupled Lattice Bisemirings,
  • from the geometric side, the category of points of the Topos $\widehat{\mathbb{N}^{\times}}$ results equivalent to several algebraic categories,
  • from the logical side we have that the theory of points of $\widehat{\mathbb{N}^{\times}}$ is recursively axiomatizable, and the theory of flat actions is geometric.
As a final comment we can say that the Tropicalization of Universal Geometry of $\L$ukasiewicz logics ([35], [3]) seems to be a rich topic which deserves to be addressed. This means extending the perspective of the interactions exhibited in the present article.

### v2

@@ 4 MV-algebras and the topos $\widehat{\mathbb{N}^{\times}}$
Geometric logic is a fragment of infinitary first order logic that focuses on formulas and reasoning preserved under colimits. Unlike classical logic, which emphasizes arbitrary conjunctions and negations, geometric logic emphasizes finite conjunctions, arbitrary disjunctions, and existential quantifiers. This focus makes geometric logic particularly well-suited to categorical settings and topoi.
While geometric logic primarily allows existential quantifiers, universal quantifiers can be used if the formula being universally quantified can be equivalently rewritten using finite conjunctions and negation of atomic formulas, which geometric logic permits implicitly.
A geometric theory is a collection of axioms expressed in geometric logic, and its classifying topos is a category that serves as a universal semantic model for the theory. The connection between syntactic (logical) and semantic (categorical) perspectives exemplifies the deep interplay between geometric logic and topoi.
@@
$G$ is a subgroup of $\mathbb{Q}$ if and only if $G$ is a group, $G$ is torsion free, and for every $0\neq x,y\in G$ there are $m,n\in\mathbb{N}$ such that $mx=ny$. Furthermore, denoting the order of $G$ by $<_{G}$, the following conditions hold: $\forall x$ $\forall y$ $\exists n,m\in\mathbb{Z}$ with
@@
The class $Perf^{\mathbb{Q}}$ is geometric in the language of $MV$-algebras.
@@
$\forall x$ $\forall y$ $\exists n$ $\exists m$ such that: $m,n\in{\mathbb{Z}}^{+}$
@@
and $\Theta_{pt}$ is a functor from $Pt(\widehat{\mathbb{N}^{\times}})$ to $\ell BS$.
@@
There is a natural functor from the category $Pt(\widehat{\mathbb{N}^{\times}})$ to the category of $\ell$-Bisemirings.
@@
Indeed, we have a threefold interaction among the geometric side, points of the topos $\widehat{\mathbb{N}^{\times}}$, the algebraic side of ordered subgroups of $\mathbb{Q}$ and the logical side of $Perf^{\mathbb{Q}}$, models of an extension of Łukasiewicz logic.
@@ 4.1 Flatness and $Pt(\widehat{\mathbb{N}^{\times}})$
    $F(*)=X\not=\emptyset$, where * is the unique object of the category $\mathbb{N}^{\times}$
@@ 5 Conclusions
As stated in the Introduction, the purpose of this article is to illustrate the interactions between algebraic structures such as Semirings, Tropical structures and Łukasiewicz logic.
This perspective has led to a generalisation of the concept of Tropicalisation of algebraic structures in a form that opened up an unexpected interpretation of an extension of Łukasiewicz logic. Idempotent semifields are the basic algebraic structures needed to built a tropical version of algebraic geometry: here we showed that if we focus on particular on semifields arising from subclasses of MV-algebras, we can connect the tropical geometry with the rich world of logic, through a topos prespective.
Considering the functorial version of the tropicalization of an algebraic structure, the above interation, among others, allows to show that from the geometric side, the category of points of the topos $\widehat{\mathbb{N}^{\times}}$ results equivalent to several algebraic categories.
As possible continuation of this work, the topics in Section 4 can be extended to find computability and complexity results of first order theories of MV-algebras and of points of the topos $\mathbb{N}^{\times}$. Further, the connections between the logic of perfect MV-algebras, topos and categories arising from idempotent semirings still need a deep investigation.